OpenStudy (anonymous):

Simplify the expression: 2x^2 y 9xy^3 over times over 3x 4y Simplify the expression: 2x^2 y 9xy^3 over times over 3x 4y @Mathematics

OpenStudy (anonymous):

in what grade u r @ selena

OpenStudy (anonymous):

Junior , but i had to retake algebra 1

OpenStudy (anonymous):

y.....are u poor @ algebra

OpenStudy (anonymous):

Yeah, it's not one of my best subjects.

OpenStudy (anonymous):

\[\frac{2x^{2}y}{3x} \times \frac{9xy^{3}}{4y}\] can u try to solve this one???

jhonyy9 (jhonyy9):

2xy/3 * 9xy2/4 = 3x2y3/2